The position of a particle moving along the x axis is given by x = 5 + 6t -3t2 meters, where t is in seconds. What is the average velocity during the time interval t = 2.0s to t=4.0s?
Answer #1


dt v12.on.) G-6x2= -6 m[Acc v(u.onu ) = 6-6xy = -1B mhec -12